# Remarks on classical pseudo-electrodynamics

### v1

@@ I Introduction
The connection between condensed matter field and field theories has gained relevance in the last decades due to applications in bidimensional materials. The quantum Hall effect [1, 2, 3, 4], planar materials [5, 6, 7, 8, 9], transport in graphene [10, 11, 12, 13], and superconductivity [14, 15, 16, 17, 18] are the examples in which the quantum field theories in two dimensions show successful description of the quantum properties in material physics. The interaction of fermions in planar materials is the main motivation to understand it mediator field theory. Based on principles of the quantum electrodynamics (QED), the structure is a gauge field theory in $1+2$ dimensions whose the potential links the interaction electron-electron. This gauge theory is known in the literature as pseudo-electrodynamics (PED) [19].
The PED is a non-local electromagnetism that arises from the dimensional reduction of Maxwell’s electrodynamics when the classical sources are constrained to a two-dimensional space. As consequence, the theory dimensionally reduced is non-local with derivatives of infinity order in the D’Alembertian operator. As fundamental properties, the PED preserves the gauge invariance, the theory respects the causality principle [20], and from the point of view of a quantum field theory, the unitarity is also confirmed [21]. Some extensions of the PED has been studied in the literature in last times, as the addition of Chern-Simons topological term [22, 23], the dimensional reduction in Proca electrodynamics [24, 25, 26], and the Lee-Wick PED if the interest it is studying a massive pseudo-gauge field with gauge invariance [27].
In this work, we show some classical results of the pseudo-electrodynamics. We start with a review on the PED. From the action of PED, we calculate the energy-momentum tensor and the conserved current via Noether theorem. Since the energy-momentum tensor is not gauge invariant and also is not a symmetric tensor, we construct the approach to find the Belinfante-Rosenfeld tensor, that yields the new energy momentum tensor, symmetric and gauge invariant. Thereby, we obtain the energy, the linear momentum, the general angular momentum, and the Lorentz force law that provides the stress tensor stored in the fields of the PED. Using the retarded solutions for the potentials and Jefimenko’s equations, we apply it to the case of a time-varying electric dipole placed in a two-dimensional plane in the approximation of the radiation zone. Thus, we calculate the power radiated in relation to radial distance. Posteriorly, we present a way to introduce magnetic charge and current distributions in the PED.
@@ II A review of pseudo-electrodynamics
where ${\cal M}$ symbolizes the Minkowski spacetime $1+3$, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}=(-E^{i},-\epsilon^{ijk}\,B^{k})$ is the electromagnetic field tensor, $A^{\mu}$ is the four-potential, $J^{\mu}$ are the classical sources, and $\xi$ is a real gauge parameter. As usual, the action principle yields the field equation for the $A^{\mu}$-potential
@@
Using the Fourier integral, the result for $\Delta^{\mu\nu}(x-x^{\prime})$ is
@@
and substituting (3), the action is a functional of $J^{\mu}$
@@
We have discarded the gauge term of $k_{\bar{\mu}}\,k_{\bar{\nu}}$ by the condition of the conserved current $k_{\bar{\mu}}\,j^{\bar{\mu}}=0$ in the momentum space when it is inserted in the action (9). Here, $j^{\bar{\mu}}=(\sigma,K^{i})$ sets a superficial current density. Since $\bar{k}^{2}=k_{0}^{2}-k_{x}^{2}-k_{y}^{2}>0$, or $\bar{k}^{2}<0$, it is convenient to introduce the coordinates in the Euclidean space-time, where $x_{0}=-i\,x_{4}$ and $k_{0}=i\,k_{4}$, such that $\bar{k}^{2}=-(k_{4}^{2}+k_{x}^{2}+k_{y}^{2})=-\bar{k}_{E}^{2}<0$. Thereby, the Green function in Euclidean (and reduced) space-time is
@@
where $F_{\bar{\mu}\bar{\nu}}=\partial_{\bar{\mu}}A_{\bar{\nu}}-\partial_{\bar{\nu}}A_{\bar{\mu}}=(-E^{i},-B_{z})\,(i=1,2)$ is the EM strength field tensor, $A^{\bar{\mu}}=(A^{0},A^{i})$ is the $3$-potential, the operator with bar index means the derivatives in relation to $(t,x,y)$, $\partial_{\bar{\mu}}=(\partial_{t},\partial_{x},\partial_{y})$, and ¯$\bar{\Box}=\partial_{\bar{\mu}}\partial^{\bar{\mu}}=\partial_{t}^{2}-\partial_{x}^{2}-\partial_{y}^{2}$ is the D’Alembertian operator in $1+2$ dimensions. The non-local operator $2/\sqrt{\bar{\Box}}$ in the kinetic term can be written as the infinite series
@@
and the Bianchi identity keeps like in the usual ED in $1+2$ dimensions
@@
where $t_{r}=t-|\bar{{\bf r}}-\bar{{\bf r}}^{\prime}|$ is the retarded time, and the integral is over the spatial plane $\mathcal{S}$. Thereby, the non-local operator of this theory does not change the interpretation of the retarded propagation for any source of charges and currents on the spatial plane, and as consequence, the Green function maintains the causality in the PED [28].
@@ III Noether theorem and the Belinfante-Rosenfeld tensor in PED
The concepts of symmetries and conserved quantities are linked by the Noether theorem. The invariance of the action under infinitesimal transformations of the space-time coordinates and of fields leads to physical conserved quantities. We start with the infinitesimal transformations in the coordinates, and the in $3$-potential of the PED
@@
where $\delta x^{\bar{\mu}}$ is a small variation in the coordinates $x^{\bar{\mu}}$. The transformations (22a) and (22b) lead to variation in the lagrangian of PED as, $\mathcal{L}_{PED}\,\longmapsto\,\mathcal{L}^{\prime}_{PED}=\mathcal{L}_{PED}+\delta\mathcal{L}_{PED}$, and the action variation is
@@
The variation of $A^{\bar{\mu}}$ in the same point $x^{\bar{\mu}}$ is defined by $\tilde{\delta}A^{\bar{\nu}}=A^{{}^{\prime}\bar{\nu}}(x)-A^{\bar{\nu}}(x)$, that can be written as $\tilde{\delta}A^{\bar{\nu}}=\delta A^{\bar{\nu}}-(\partial_{\bar{\mu}}A^{\bar{\nu}})\;\delta x^{\bar{\mu}}$. The variation of the Lagrangian in $x^{\bar{\mu}}$ is $\tilde{\delta}\mathcal{L}_{PED}=\delta\mathcal{L}_{PED}-(\partial_{\bar{\mu}}\mathcal{L}_{PED})\,\delta x^{\bar{\mu}}$. Using these results in (III), the variation of the action is
@@
where $\omega^{\bar{\alpha}\bar{\beta}}$ are real parameters, in which $\omega^{\bar{\alpha}\bar{\beta}}=-\omega^{\bar{\beta}\bar{\alpha}}$, and $(\Sigma^{\bar{\alpha}\bar{\beta}})_{\bar{\nu}\bar{\sigma}}=i\left(\delta^{\bar{\alpha}}_{\;\;\,\bar{\sigma}}\,\delta^{\bar{\beta}}_{\;\;\,\bar{\nu}}-\delta^{\bar{\alpha}}_{\;\;\,\bar{\nu}}\,\delta^{\bar{\beta}}_{\;\;\,\bar{\sigma}}\right)$ is the generator for the vector representation in the Lorentz group. Substituting these expressions in (27), the Noether current is written as
@@
in which $N^{\bar{\nu}\bar{\alpha}}=-N^{\bar{\alpha}\bar{\nu}}$ is antisymmetric, and therefore, it has three components $N^{\bar{\nu}\bar{\alpha}}=(N^{01},N^{02},N^{12})$ in $1+2$ dimensions. The spatial component is defined as the general angular momentum on $Z$-direction, i.e., $N^{12}=J_{z}=L_{z}+S_{z}$, where the orbital angular momentum $(L_{z})$ and the spin $(S_{z})$ are given by
@@
that satisfies the properties of permutation of the index, shown previously, in which $Y_{\bar{\sigma}\bar{\mu}\bar{\nu}\bar{\lambda}}=x_{\bar{\nu}}\,X_{\bar{\sigma}\bar{\mu}\bar{\lambda}}-x_{\bar{\lambda}}\,X_{\bar{\sigma}\bar{\mu}\bar{\nu}}$. Therefore, the full energy-momentum symmetric tensor is
@@
The conservation laws for the symmetric tensor $T^{\bar{\mu}\bar{\nu}}$ also can be obtained directly from the field equations (17) and (18). Contracting (17) with $F_{\bar{\nu}\bar{\rho}}$, with help of (18), we obtain
@@
This result is usefull for the calculus of force and pressure on the surfaces due to external sources, as example. Thereby, we have obtained all the conserved physical quantities that are gauge invariant in the PED. These quantities can yields results of physical interest as the energy and the radiation power that will be investigated in the next section.
@@ IV Classical radiation in pseudo-electrodynamics
The Green’s function method shows that the solution for the $A^{\bar{\mu}}$-potential in PED is similar to the Maxwell ED. Thus, the causality principle is respected in PED even if it is a non-local theory. The retarded solutions for the components of the potential are read as
@@ V Pseudo-electrodynamics with magnetic monopoles
From PED equations, there is not the divergent of the magnetic field in $1+2$ dimensions, and therefore, the introduction of a magnetic charge density is not direct like in the case of the Maxwell ED in $1+3$ dimensions. In the context of the PED, we introduce the new strength field tensor associated with the electric charge as $F_{\bar{\mu}\bar{\nu}}\rightarrow F_{e\bar{\mu}\bar{\nu}}=\partial_{\bar{\mu}}A_{e\bar{\nu}}-\partial_{\bar{\nu}}A_{e\bar{\mu}}=(-E_{e}^{\,\,i},B_{ez})$, and also, the strength field tensor associated with the magnetic charge distribution, that is, $F_{m\bar{\mu}\bar{\nu}}=\partial_{\bar{\mu}}A_{m\bar{\nu}}-\partial_{\bar{\nu}}A_{m\bar{\mu}}=(-B_{m}^{\,\,i},E_{mz})$, where $A_{m}^{\,\,\bar{\mu}}=(\Phi_{m},{\bf A}_{m})$ is the magnetic $3$-potential, and $A_{e}^{\,\,\bar{\mu}}=(\Phi_{e},{\bf A}_{e})$ is the electric $3$-potential. Based on the approach for the usual ED in $1+2$ dimensions [30], the PED in the presence of magnetic charges and currents is represented by the set of equations
@@
in which $j_{m}^{\,\,\bar{\nu}}=(\sigma_{m},{\bf K}_{m})$ is the magnetic current on the spatial plane, and $j_{e}^{\,\,\bar{\nu}}=(\sigma_{e},{\bf K}_{e})$ is the electric current. The matrix representation for EM tensors is
@@
Equations (63a) and (63b) are invariant under the duality symmetry $(E_{e}^{\,\,i},B_{ez})\rightarrow(B_{m}^{\,\,i},-E_{mz})$, $(\sigma_{e},{\bf K}_{e})\rightarrow(\sigma_{m},{\bf K}_{m})$ and $(B_{m}^{i},E_{mz})\rightarrow(-E_{e}^{\,\,i},B_{ez})$, $(\sigma_{m},{\bf K}_{m})\rightarrow(-\sigma_{e},-{\bf K}_{e})$. The gauge symmetry is set by the transformations of the electric and magnetic potentials, i.e., $A^{\prime}_{e\bar{\mu}}=A_{e\bar{\mu}}-\partial_{\bar{\mu}}\Lambda_{e}$ and $A^{\prime}_{m\bar{\mu}}=A_{m\bar{\mu}}-\partial_{\bar{\mu}}\Lambda_{m}$, for any two scalar functions $\Lambda_{e}$ and $\Lambda_{m}$, respectively. The total EM field is defined by $F_{\bar{\mu}\bar{\nu}}=F_{e\bar{\mu}\bar{\nu}}+F_{m\bar{\mu}\bar{\nu}}$ , that obviously is gauge invariant. The variational principle in relation to the potentials $A_{e\bar{\mu}}$ and $A_{m\bar{\mu}}$ (that are treated as independent) is insured by the lagrangian density
@@
in which $t_{r}=t-|{\bf r}-{\bf r}^{\prime}|$ is the retarded time as defined previously.
The conservation laws can be obtained from the field equations (63a) and (63b), using the similar approach from section (III). Multiplying the source equations by $F_{e\bar{\nu}\bar{\rho}}$ and $F_{m\bar{\nu}\bar{\rho}}$, respectively, the sum of these equations with the Bianchi identities yields the continuity equation in the presence of external sources
@@
This expression contains the result (41) if we set $B_{m}^{\,i}=E_{mz}=0$. Thus, the structure of the PED with magnetic monopoles is constructed through the new gauge potential that contains all the information of charge/current for magnetic distributions, and in principle, are independent on the equations from the electric sector.

### v2

@@ I Introduction
The connection between condensed matter field and field theories has gained relevance in the last decades due to applications in bidimensional materials. The quantum Hall effect [1, 2, 3, 4], planar materials [5, 6, 7, 8, 9], transport in graphene [10, 11, 12, 13], and superconductivity [14, 15, 16, 17, 18] are the examples in which the quantum field theories in two dimensions show a successful description of the quantum properties in material physics. The interaction between fermions in planar materials is the main motivation to understand the mediator field theory. Based on principles of the quantum electrodynamics (QED), the structure is a gauge field theory in $1+2$ dimensions whose the potential links the interaction electron-electron. This gauge theory is known in the literature as pseudo-electrodynamics (PED) [19].
The PED is a non-local electromagnetism that arises from the dimensional reduction of the Maxwell’s electrodynamics when the classical sources are constrained to a two-dimensional space. As consequence, the theory dimensionally reduced is non-local with derivatives of infinity order in the D’Alembertian operator. As fundamental properties, the PED preserves the gauge invariance, the theory respects the causality principle [20], and from the point of view of a quantum field theory, the unitarity is also confirmed [21]. Some extensions of the PED has been studied in the literature in last times, as the addition of Chern-Simons topological term [22, 23], the dimensional reduction in Proca electrodynamics [24, 25, 26], and the Lee-Wick PED, if the interest it is studying a massive pseudo-gauge field with gauge invariance [27]. The introduction of magnetic monopoles in PED also is motivated in two-dimensional topological insulators and in superconductors [28]. The formulation of two-dimensional magnetic monopole gas is another way to investigate the emergence of magnetic monopoles in semiconductor heterostructures [29].
In this work, we show some classical results of the pseudo-electrodynamics. We start with a review on the PED. From the action of the PED, we calculate the energy-momentum tensor and the conserved current via Noether theorem. Since the energy-momentum tensor is not gauge invariant, and it is not a symmetric tensor, we construct the approach to find the Belinfante-Rosenfeld tensor, that yields the new energy momentum tensor, symmetric and gauge invariant. Thereby, we obtain the energy, the linear momentum, the general angular momentum, and the Lorentz force law that provides the stress tensor stored in the fields of the PED. Using the retarded solutions for the potentials and Jefimenko’s equations, we apply it to the case of a time-varying electric dipole placed in a two-dimensional plane in the approximation of the radiation zone. Thus, we calculate the power radiated in relation to radial distance. Posteriorly, we present a way to introduce magnetic charge and current distributions in the PED.
@@ II A review of pseudo-electrodynamics
where ${\cal M}$ symbolizes the Minkowski spacetime $1+3$, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}=(-E^{i},-\epsilon^{ijk}\,B^{k})$ is the electromagnetic field tensor, $A^{\mu}$ is the four-potential, $J^{\mu}$ are the classical sources, and $\xi$ is a gauge fixing parameter. As usual, the action principle yields the field equation for the $A^{\mu}$-potential
@@
Using the Fourier integral, the result for $\Delta_{\mu\nu}(x-x^{\prime})$ is
@@
and substituting (3), the action is now a functional of $J^{\mu}$
@@
We have discarded the gauge term of $k_{\bar{\mu}}\,k_{\bar{\nu}}$ by the condition of the conserved current $k_{\bar{\mu}}\,j^{\bar{\mu}}=0$ in the momentum space when the Green function is inserted in the action (9). Here, $j^{\bar{\mu}}=(\sigma,K^{i})$ sets a superficial current density. Since $\bar{k}^{2}=k_{0}^{2}-k_{x}^{2}-k_{y}^{2}>0$, or $\bar{k}^{2}<0$, it is convenient to introduce the coordinates in the Euclidean space-time, where $x_{0}=-i\,x_{4}$ and $k_{0}=i\,k_{4}$, such that $\bar{k}^{2}=-(k_{4}^{2}+k_{x}^{2}+k_{y}^{2})=-\bar{k}_{E}^{2}<0$. Thereby, the Green function in Euclidean (and reduced) space-time is
@@
where $F_{\bar{\mu}\bar{\nu}}=\partial_{\bar{\mu}}A_{\bar{\nu}}-\partial_{\bar{\nu}}A_{\bar{\mu}}=(-E^{i},-B_{z})\,(i=1,2)$ is the EM strength field tensor, $A^{\bar{\mu}}=(A^{0},A^{i})$ is the $3$-potential, the operator with bar index means the derivatives in relation to $(t,x,y)$, $\partial_{\bar{\mu}}=(\partial_{t},\partial_{x},\partial_{y})$, and ¯$\bar{\Box}=\partial_{\bar{\mu}}\partial^{\bar{\mu}}=\partial_{t}^{2}-\partial_{x}^{2}-\partial_{y}^{2}$ is the D’Alembertian operator in $1+2$ dimensions. The non-local operator $2/\sqrt{\bar{\Box}}$ in the kinetic term can be written as the infinity series
@@
and the Bianchi identity keeps like in the usual ED, but with the bar index in $1+2$ dimensions
@@
where $t_{r}=t-|\bar{{\bf r}}-\bar{{\bf r}}^{\prime}|$ is the retarded time, and the integral is over the spatial plane $\mathcal{S}$. Thereby, the non-local operator of this theory does not change the interpretation of the retarded propagation for any source of charges and currents on the spatial plane, and as consequence, the Green function preserves the causality in the PED [30].
@@ III Noether theorem and the Belinfante-Rosenfeld tensor in PED
The concepts of symmetries and conserved quantities are linked by the Noether theorem. The invariance of the action under infinitesimal transformations of the space-time coordinates and of fields leads to physical conserved quantities. We start with the infinitesimal transformations in the coordinates, and in the $3$-potential of the PED
@@
where $\delta x^{\bar{\mu}}$ is a small variation of the coordinates $x^{\bar{\mu}}$. The transformations (22a) and (22b) lead to variation in the lagrangian of the PED as, $\mathcal{L}_{PED}\,\longmapsto\,\mathcal{L}^{\prime}_{PED}=\mathcal{L}_{PED}+\delta\mathcal{L}_{PED}$, and the action variation is
@@
The variation of $A^{\bar{\mu}}$ in the same point $x^{\bar{\mu}}$ is defined by $\tilde{\delta}A^{\bar{\nu}}=A^{{}^{\prime}\bar{\nu}}(x)-A^{\bar{\nu}}(x)$, that can be written as $\tilde{\delta}A^{\bar{\nu}}=\delta A^{\bar{\nu}}-(\partial_{\bar{\mu}}A^{\bar{\nu}})\;\delta x^{\bar{\mu}}$. Thereby, the variation of the Lagrangian in $x^{\bar{\mu}}$ is $\tilde{\delta}\mathcal{L}_{PED}=\delta\mathcal{L}_{PED}-(\partial_{\bar{\mu}}\mathcal{L}_{PED})\,\delta x^{\bar{\mu}}$. Using these results in (III), the variation of the action is
@@
where $\omega^{\bar{\alpha}\bar{\beta}}$ are real parameters, in which $\omega^{\bar{\alpha}\bar{\beta}}=-\omega^{\bar{\beta}\bar{\alpha}}$, and $(\Sigma^{\bar{\alpha}\bar{\beta}})_{\bar{\nu}\bar{\sigma}}=i\left(\delta^{\bar{\alpha}}_{\;\;\,\bar{\sigma}}\,\delta^{\bar{\beta}}_{\;\;\,\bar{\nu}}-\delta^{\bar{\alpha}}_{\;\;\,\bar{\nu}}\,\delta^{\bar{\beta}}_{\;\;\,\bar{\sigma}}\right)$ is the generator for the vector representation in the Lorentz group. Substituting these variations in (27), the Noether current is written as
@@
in which $N^{\bar{\nu}\bar{\alpha}}=-N^{\bar{\alpha}\bar{\nu}}$ is antisymmetric, and therefore, it has three components $N^{\bar{\nu}\bar{\alpha}}=(N^{01},N^{02},N^{12})$ in $1+2$ dimensions. The spatial component is defined as the general angular momentum on ${\cal Z}$-direction, i.e., $N^{12}=J_{z}=L_{z}+S_{z}$, where the orbital angular momentum $(L_{z})$ and the spin $(S_{z})$ are given by
@@
that satisfies the properties of permutation of the index, shown previously, in which $Y_{\bar{\sigma}\bar{\mu}\bar{\nu}\bar{\lambda}}=x_{\bar{\nu}}\,X_{\bar{\sigma}\bar{\mu}\bar{\lambda}}-x_{\bar{\lambda}}\,X_{\bar{\sigma}\bar{\mu}\bar{\nu}}$. Therefore, the symmetric energy-momentum tensor is
@@
Alternatively, the conservation laws for the symmetric tensor $T^{\bar{\mu}\bar{\nu}}$ also can be obtained directly from the field equations (17) and (18). Contracting (17) with $F_{\bar{\nu}\bar{\rho}}$, with help of (18), we obtain
@@
This result is usefull for the calculus of force and pressure on the surfaces due to external sources, as example. Thereby, we have obtained all the conserved physical quantities that are gauge invariant in the PED. These quantities can yield results of physical interest as the energy and the radiation power that will be investigated in the next section.
@@ IV Classical radiation in pseudo-electrodynamics
The Green’s function method shows that the solution for the $A^{\bar{\mu}}$-potential in PED is similar to the Maxwell ED. Thus, the causality principle is preserved in PED, even if it is a non-local theory [30]. The retarded solutions for the components of the potential are read as
@@ V Pseudo-electrodynamics with magnetic monopoles
From the PED equations, there is not the divergent of the magnetic field in $1+2$ dimensions, and therefore, the introduction of a magnetic charge density is not direct like in the case of the Maxwell ED in $1+3$ dimensions. In the context of the PED, we introduce the new strength field tensor associated with the electric charge as $F_{\bar{\mu}\bar{\nu}}\rightarrow F_{e\bar{\mu}\bar{\nu}}=\partial_{\bar{\mu}}A_{e\bar{\nu}}-\partial_{\bar{\nu}}A_{e\bar{\mu}}=(-E_{e}^{\,\,i},B_{ez})$, and also, the strength field tensor associated with the magnetic charge distribution, that is, $F_{m\bar{\mu}\bar{\nu}}=\partial_{\bar{\mu}}A_{m\bar{\nu}}-\partial_{\bar{\nu}}A_{m\bar{\mu}}=(-B_{m}^{\,\,i},E_{mz})$, where $A_{m}^{\,\,\bar{\mu}}=(\Phi_{m},{\bf A}_{m})$ is the magnetic $3$-potential, and $A_{e}^{\,\,\bar{\mu}}=(\Phi_{e},{\bf A}_{e})$ is the electric $3$-potential. Based on the approach for the usual ED in $1+2$ dimensions [32], the PED in the presence of magnetic charges and currents is represented by the set of equations
@@
in which $j_{m}^{\,\,\bar{\nu}}=(\sigma_{m},{\bf K}_{m})$ is the magnetic $3$-current on the spatial plane, and $j_{e}^{\,\,\bar{\nu}}=(\sigma_{e},{\bf K}_{e})$ is the electric $3$-current. The matrix representation for EM tensors is
@@
Equations (63a) and (63b) are invariant under the duality symmetry $(E_{e}^{\,\,i},B_{ez})\rightarrow(B_{m}^{\,\,i},-E_{mz})$, $(\sigma_{e},{\bf K}_{e})\rightarrow(\sigma_{m},{\bf K}_{m})$ and $(B_{m}^{i},E_{mz})\rightarrow(-E_{e}^{\,\,i},B_{ez})$, $(\sigma_{m},{\bf K}_{m})\rightarrow(-\sigma_{e},-{\bf K}_{e})$. The gauge symmetry is represented by the transformations of the electric and magnetic potentials, i.e., $A^{\prime}_{e\bar{\mu}}=A_{e\bar{\mu}}-\partial_{\bar{\mu}}\Lambda_{e}$ and $A^{\prime}_{m\bar{\mu}}=A_{m\bar{\mu}}-\partial_{\bar{\mu}}\Lambda_{m}$, for any two scalar functions $\Lambda_{e}$ and $\Lambda_{m}$, respectively. The total EM field is defined by $F_{\bar{\mu}\bar{\nu}}=F_{e\bar{\mu}\bar{\nu}}+F_{m\bar{\mu}\bar{\nu}}$ , that obviously is gauge invariant. The variational principle in relation to the potentials $A_{e\bar{\mu}}$ and $A_{m\bar{\mu}}$ (that are treated as independent) is insured by the lagrangian density
@@
in which $t_{r}=t-|{\bf r}-{\bf r}^{\prime}|$ is the retarded time, as defined previously.
The conservation laws can be obtained from the field equations (63a) and (63b), using the similar approach from the section (III). Multiplying the source equations by $F_{e\bar{\nu}\bar{\rho}}$ and $F_{m\bar{\nu}\bar{\rho}}$, respectively, the sum of these equations with the Bianchi identities yields the continuity equation in the presence of external sources
@@
This expression contains the result (41) if we set $B_{m}^{\,i}=E_{mz}=0$. Thus, the structure of the PED with magnetic monopoles is constructed through the new gauge potential that contains all the information of charge/current for magnetic distributions, and in principle, are independent on the equations of the electric sector. The result (77) can lead to a new quantization rule for the magnetic monopole in the pseudo-electrodynamics, that will be investigated in a forthcoming project.